Instant hot packs work by crystallizing sodium acetate (NaCH3COO). The molar enthalpy of crystallization for sodium acetate is -56.7 kJ/mol. How many grams of sodium acetate are needed to warm 125.0 ml of water from 21 degrees C to 35.4 degrees C?


1
Expert's answer
2021-01-05T05:56:32-0500

"V = 125.0ml"

Since, "m = \rho × V = 1g/ml × 125ml = 125g"

"\u2206T = T_2 - T_1 = 35.4\u00b0C - 21\u00b0C = 14.4\u00b0C"

"c = 4.186J\/g\u00b0C"


H = mc∆T

H = 125 × 4.186 × 14.4 = 7534.8J = 7.535kJ



Since the molar enthalpy of crystallization for sodium acetate is -56.7 kJ/mol

56.7kJ = 1 mole of sodium acetate

7.535kJ = x moles


x = "\dfrac{7.535×1}{56.7}" = 0.133 moles



1 mole of "CH_3COONa" = 82g

0.133 moles = xg


x = 0.133 × 82 = 10.9g



This means that 10.9 grams of sodium acetate is needed to warm 125.0 ml of water from 21°C to 35.4°C.